Simplify -2√6 -2√36 -4√54
1
Expert's answer
2012-02-07T09:44:10-0500
-2 sqrt(6)-2 sqrt(36)-4 sqrt(54)=
-2 sqrt(6)-2 sqrt(6*6)-4 sqrt(6*9)=
-2 sqrt(6)-2*6-4*3 sqrt(6)=-14 sqrt(6)-12